A rower travels upstream at 6 km per hour and back to the starting place at 10 km per hour. the total journey takes 48 minutes. How far upstream did the rower go?

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

4/5=48/60
a+b=4/5,
6a=8b
a = 16/35, b = 12/35
96/35 km =distance
2 26/35 k=distance
2.7428571 k=distance

Answer by Alan3354(69443)

A rower travels upstream at 6 km per hour and back to the starting place at 10 km per hour. the total journey takes 48 minutes. How far upstream did the rower go?
============
Avg speed for the round trip is
2*6*10/(6+10) = 7.5 km/hr
d = r*t
d = 7.5*48/60) = 6 km round trip
--> 3 km each way
-------------------
I can show the derivation of the RT avg speed if necessary.